A force of 6.4 N acts on a 22 kg body initially at rest. Compute the work done by the force in (a) the first, (b) the second, and (c) the third seconds and (d) the instantaneous power due to the force at the end of the third second.
Answer #1

hiven datur force, F = 6.4 N mass, n=22 kg Instal Velocity Vi= o mas acceleration, a=t= 6-4 = 0.2909 mis 22 distance levered,☺ one Sin the teamed W3 = f ( t at ) - F ( 2 at 3) - Flitats) - (W2 twi) = 6-4 x4 x0-2969432- (093+2-8) | W3 = 4.645) Instan